OpenStudy (anonymous):

Quadrilateral ABCD with vertices at A(-5,-1),B (2,4),C (9, 2), and D (2,-3) is shown. Diagonal BD is drawn. Use coordinate geometry to prove that angels ABD=CDB.

OpenStudy (anonymous):



OpenStudy (anonymous):

...it looks like that

OpenStudy (anonymous):

Good drawing!

OpenStudy (anonymous):

The first thing is your points need to have two numbers each. Coordinate pairs have an x value and a y value (x,y).

OpenStudy (anonymous):

well ya they do in the question :)

OpenStudy (anonymous):

Right, your drawing should have that information too. What are the numbers in your drawing?

OpenStudy (anonymous):

Ahem, diagram*.

OpenStudy (anonymous):



OpenStudy (anonymous):

Good.

OpenStudy (anonymous):

-5-1 lol I messed up on the picture xp

OpenStudy (anonymous):

That's okay, just remember it's (-5, -1) for the calculation.

OpenStudy (anonymous):

Okay, next question, what is the distance formula?

OpenStudy (anonymous):

lol I thought you were just going to give me the answer so its not long and confusing XD

OpenStudy (anonymous):

Use the distance formula to find the length of AB, BC, BD, CD, and AD.

OpenStudy (anonymous):

Distance Formula \[distance = \sqrt{(x _{1} - x _{2})^{2} + (y _{1} - y _{2})^{2}}\]

OpenStudy (anonymous):

Lets find AB together using the distance formula.

OpenStudy (anonymous):

3.4?

OpenStudy (anonymous):

8.6!?

OpenStudy (anonymous):

\[AB = \sqrt{(-5 - 2)^{2} + (-1 - 4)^{2}}\] \[AB = \sqrt{(-7)^{2} + (-5)^{2}}\] \[AB = \sqrt{49 + 25}\] \[AB = \sqrt{74}\]

OpenStudy (anonymous):

ya 8.6 lol

OpenStudy (anonymous):

Great!

OpenStudy (anonymous):

Okay, now use the distance formula to find BC, CD, AD, and BD.

OpenStudy (anonymous):

ok now the whoe final answer all together is what lmao =p

OpenStudy (anonymous):

whole

OpenStudy (anonymous):

This is a proof problem that requires multiple steps. Thank you for being so patient with this.

OpenStudy (anonymous):

Okay, find BC using the distance formula.

OpenStudy (anonymous):

lol 7.2

OpenStudy (anonymous):

BC = sqrt( (2 - 9)^2 + (4 - 2)^2) BC = sqrt( (-7)^2 + (-2)^2) BC = sqrt( 49 + 4) BC = sqrt(54) BC = 3 * sqrt(6)

OpenStudy (anonymous):

hmmm, I'm getting 7.34.

OpenStudy (anonymous):

Can you show me any of the work that you are doing to get your answers?

OpenStudy (anonymous):

What was the number under the square root?

OpenStudy (anonymous):

Ah, I found my mistake.